OpenStudy (anonymous):

Differentiate: sqrt{2+t^2} u = 2 + t^2 dt = du/2t int_{}{}u ^{1/2} dt (u ^{3/2})/(3/2) * du/2t Answer: (2+t^2)^{3/2} / 3t Is it correct?

OpenStudy (anonymous):

\[2t*1/(\sqrt{2+t^2})\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{2+t^2}=(2+t^2)^{\frac{1}{2}}\] \[\Large \frac{d}{dx}\left(\sqrt{2+t^2}\right)=\frac{d}{dx}\left((2+t^2)^{\frac{1}{2}}\right)\] \[\Large \frac{d}{dx}\left(\sqrt{2+t^2}\right)=\frac{1}{2}(2+t^2)^{-\frac{1}{2}}\times \frac{d}{dx}(2+t^2)\] \[\Large \frac{d}{dx}\left(\sqrt{2+t^2}\right)=\frac{2t}{2\sqrt{2+t^2}}\] \[\Large \frac{d}{dx}\left(\sqrt{2+t^2}\right)=\frac{t}{\sqrt{2+t^2}}\]

OpenStudy (anonymous):

Corrected formatting: \[\sqrt{2+t^2}\] \[u = 2 + t^2\] \[dt = du/2t\] \[\int\limits\limits_{}{}u ^{1/2} dt\] \[(u ^{3/2})/(3/2) * du/2t \] \[Answer: (2+t^2)^{3/2} / 3t\]

OpenStudy (anonymous):

oops